OpenStudy (anonymous):

What does it mean when it says substitute 0 for the function notation? can you give me an example of an equation that is substituted by 0 please?

OpenStudy (anonymous):

I would guess that it means that you have some function f(x) = blah and you are to put 0 = blah in order to find roots or something of that sort...

OpenStudy (anonymous):

it asked me to do this and then it asks me that function question, 5. Rational Root system: I checked for roots at x = ± (1, 2, 3, 4, 6, 9, 12, 18, 27, 36, 54, and 108). Though since Descartes already told me there are no negative roots, I only checked the positive roots. There is a root at x = 9. Using the factor theorem I converted it to: (x - 9) & divided the original function by the known factor: (x^3-10x^2+21x -108) / (x - 9) = x^2 - x + 12 Using the quadratic formula: x = (1/2) (1 ± i √47) The three roots of x^3-10x^2+21x -108 = 0 are: x = 9, (1/2) (1 + i √47), and (1/2) (1 - i √47)

OpenStudy (anonymous):

What I said before... I suppose you have f(x) = x^3-10x^2+21x -108, yes? So put x^3-10x^2+21x -108 = 0 and solve?

OpenStudy (anonymous):

That doesnt work bc im supposed to graph it and it just comes out as a straight line, its supposed to be a polynomial

OpenStudy (anonymous):

x^3-10x^2+21x -108 is not a straight line.

OpenStudy (anonymous):

im supposed to use geobegra and i type that in and i get a line

OpenStudy (anonymous):

I suppose you will get a line but it will be curved....

OpenStudy (anonymous):

why wont it do that in geogebra though, i need to copy and paste it from their

OpenStudy (anonymous):

No idea, sorry (I don't use geogebra)

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Hope you can sort it out.....